How many molecules are present in 500 milimole of AlCl3? How many atoms is this.

 






1
Expert's answer
2021-02-15T03:22:38-0500

According to the Avogadro's number, one mole of any substance contains 6.022 × 10²³ particles. From here, 500 mmol of AlCl3 contains:

n(AlCl3) = 6.022 × 10²³ × 500 mmol = 6.022 × 10²³ × 0.5 mol = 3.011 × 10²³ molecules

As one molecule of AlCl3 consists off four atoms, the number of atoms is 500 mmol of AlCl3 is:

n(atoms) = 3.011 × 10²³ × 4 = 1.2044 × 10²⁴


Answer: 6.022 × 10²³ molecules; 1.2044 × 10²⁴ atoms